the length of a rectangle is 12 cm greater than its width, the perimeter of this rectangle is 60 cm. find its dimension.
Answer by Cromlix(4381)

Width = x
Length = 12 + x
Perimeter = 2*width + 2*length
60 = 2(x) + 2(12 + x)
60 = 2x + 24 + 2x
4x = 60 - 24
4x = 36
x = 9cm
Width = 9cm
Length = 21cm.
Hope this helps.
:-)